OpenStudy (anonymous):

The square root of 200a^6b^7

OpenStudy (saifoo.khan):

\[\Large \sqrt{200a^6b^7} \to \sqrt{ 10(10)(2) a^6 b^7} \to 10a^3b^3 \sqrt{2 b}\]

OpenStudy (anonymous):

i dont get the 10 10 2 part..wouldnt the 200 turn into \[5\sqrt{8} * \sqrt{a ^{6}} * \sqrt{b ^{7}}\]

OpenStudy (saifoo.khan):